OpenStudy (anonymous):

Write an equation for the conic section : Parabola with the vertex at (1,-2) and a focus at (1,1) ???

OpenStudy (kohai):

 So it's going to roughly look like this

OpenStudy (kohai):

The equation you're going to use is (y-h)^2=4p(x-k)

OpenStudy (anonymous):

well thanks

OpenStudy (kohai):

I wasn't done...

OpenStudy (anonymous):

ok

OpenStudy (kohai):

The vertex is (h, k). So what is the value for h and what is the value for k? Hint, they've already given it to you

OpenStudy (anonymous):

1,-2

OpenStudy (kohai):

Right. (y-h)^2=4p(x-k) (h-1)^2=4p(x+2) Now, p is the distance between the vertex and the focus. So go back to the graph and count the amount of units

OpenStudy (anonymous):

ummm p=3 ??????

OpenStudy (kohai):

P is equal to 3, yes, so you multiply (4)(3)

OpenStudy (anonymous):

12

OpenStudy (kohai):

So the final equation will be?

OpenStudy (anonymous):

(y-1)^2=12(x+2) ?????

OpenStudy (anonymous):

am i right ???

OpenStudy (kohai):

Yeah I'm double checking on a graphing calculator to make sure it's right

OpenStudy (anonymous):

ok

OpenStudy (kohai):

Give me just a moment, please

OpenStudy (kohai):

Ah, I see. I was confused because graphed this equation is sideways. So, here's the actual equation, (x-1)^2=12(y+2) Just needed to switch the variables :)

OpenStudy (anonymous):

okay well tht helps alot thks :) !

OpenStudy (kohai):

You're welcome :)